# RiemannLoRA: A Unified Riemannian Framework for Ambiguity-Free LoRA Optimization

### v1

@@ 3.2 Riemannian optimization techniques
The Euclidean gradient is a direction of the fastest local increase $F$. Therefore, it is common to use the Riemannian gradient — the direction of local increase of corresponding smooth function value along the manifold, which lies in the tangent space (Absil et al., 2009, chap. 3.6). Given the Euclidean gradient $\nabla F$, one may endow the tangent space $\mathcal{T}_{X}\mathcal{M}_{r}$ with a natural scalar product and derive a formula for the direction of the local steepest ascent of $F$ alongside the manfold. This unique direction is called the Riemannian gradient and assuming that $F$ is defined on the whole $\mathbb{R}^{m\times n}$, can be computed as follows:
$\mathrm{grad}\,F(X)=P_{\mathcal{T}_{X}\mathcal{M}_{r}}\,\nabla F(X),\quad X\in\mathcal{M}_{r},$ (8)
where $P_{\mathcal{T}_{X}\mathcal{M}_{r}}$ is the orthogonal projection operator onto $\mathcal{T}_{X}\mathcal{M}_{r}$.
One may also use momentum vectors that are reprojected to the current tangent plane (vector transort, see (Absil et al., 2009, chap. 8.1)) and combine it with the Riemannian gradient at the $i$-th iteration:
$d^{(i)}=\beta\underbrace{P_{\mathcal{T}_{X^{(i)}}\mathcal{M}_{r}}\,d^{(i-1)}}_{\mathrm{momentum}}+(1-\beta)\underbrace{\mathrm{grad}\,F(X^{(i)})}_{\mathrm{gradient}}.$
Next, we want to move along the manifold in the chosen direction $\xi=-d^{(i)}\in\mathcal{T}_{X}\mathcal{M}_{r}$. This operation is commonly handled by the retraction function $R_{X}(\xi)$. The retraction, informally speaking, returns $X+\xi$ (of the rank $2r$ in general) back to the manifold of rank-$r$ matrices. In our case, we use a projection-type retraction ((Lewis and Malick, 2008, Lemma 2.1), see (Absil and Oseledets, 2015) for other variants of retraction functions):
$R_{X}(\xi)=U_{r}\Sigma_{r}V^{\top}_{r}\in\mathcal{M}_{r},$ (9)
where $U_{r}\Sigma_{r}V^{\top}_{r}$ is a rank-$r$ truncated SVD of the matrix $X+\xi$. Since $X\in\mathcal{T}_{X}\mathcal{M}_{r}$, one may obtain that $\mathrm{rank}\,(X+\xi)\leq 2r$, and the truncated SVD can be computed in low-rank format without forming dense matrices (see (6)). The asymptotic complexity of this retraction operation is $\mathcal{O}((m+n)r^{2}+r^{3})$.
Finalizing the optimization step, one may select a step size parameter $\eta^{(i)}$ and obtain the final expression:
$X^{(i+1)}=R_{X^{(i)}}\left(-\eta^{(i)}\cdot d^{(i)}\right).$
4 RiemannLoRA
4.1 Efficient computation of the Riemannian gradient
Recall that we represent $\Delta W$ from $W+\Delta W$ in two ways:
$\Delta W=A_{L}B^{\top}=AB_{R}^{\top}\in\mathcal{M}_{r}.$
Let $\mathcal{L}$ be the loss function of a model and $\eta$ be a step size, then
$\nabla_{\Delta W}\mathcal{L}(W+\Delta W)=\nabla_{W+\Delta W}\mathcal{L}\left(W+\Delta W\right)=\nabla_{Y}\mathcal{L}\left(Y\right),\quad Y=W+\Delta W,$
so according to (7) and (8), the Riemannian gradient can be derived as
$\mathrm{grad}_{\Delta W}\,\mathcal{L}(W+\Delta W)=(I-A_{L}A_{L})^{\top}\nabla_{Y}\mathcal{L}\left(Y\right)B_{R}B_{R}^{\top}+A_{L}A_{L}^{\top}\nabla_{Y}\mathcal{L}\left(Y\right).$ (10)
The calculation of a full loss gradient $\nabla_{Y}\mathcal{L}\left(Y\right)$ is computationally expensive. At the same time, we actually need to compute $\nabla_{Y}\mathcal{L}\left(Y\right)B_{R}$ and $A_{L}^{\top}\nabla_{Y}\mathcal{L}\left(Y\right)$. Using the trick from (Novikov et al., 2022), we may calculate both quantities using a single forward and backward passes with a doubled rank representation:
$\begin{split}&\nabla_{Z_{1}}\mathcal{L}\left(W+Z_{1}B_{R}^{\top}+A_{L}Z_{2}^{\top}\right)|_{Z_{1}=0,Z_{2}=B}=\nabla_{Y}\mathcal{L}(Y)B_{R},\\ &\nabla_{Z_{2}}\mathcal{L}\left(W+Z_{1}B_{R}^{\top}+A_{L}Z_{2}^{\top}\right)|_{Z_{1}=0,Z_{2}=B}=\nabla_{Y}\mathcal{L}(Y)^{\top}A_{L}.\end{split}$ (11)
Clearly, this approach can also be used to compute $\nabla_{Y}\mathcal{L}\left(Y\right)M$ and $N^{\top}\nabla_{Y}\mathcal{L}\left(Y\right)$ for arbitrary matrices $M\in\mathbb{R}^{n\times k}$ or $N\in\mathbb{R}^{m\times k}$. We will use this as a key building block for randomized SVD initialization in the next section.
4.2 Locally optimal initialization (LOI)
Given any $\Delta W\in\mathcal{M}_{r}$, we may write
$\mathcal{L}(W)=\mathcal{L}(\underbrace{W-\Delta W}_{W^{\prime}}+\Delta W)=\mathcal{L}\left(W^{\prime}+\Delta W\right).$
@@ 4.2 Locally optimal initialization (LOI)
$\begin{split}\Delta W^{(0)}_{*}\in\Bigg{\{}&\begin{bmatrix}U_{1,r},U_{r,2r}\Sigma_{r,2r}\end{bmatrix}\begin{bmatrix}S_{11}\\ S_{21}\end{bmatrix}\begin{bmatrix}C_{21}&C_{22}\end{bmatrix}\begin{bmatrix}\Sigma_{1,r}V_{1,r}^{\top}\\ V_{r,2r}\end{bmatrix}\Bigg{|}\,\\ &S=\begin{bmatrix}S_{11}&S_{12}\\ S_{21}&S_{22}\end{bmatrix}\in\mathrm{GL}_{2r}(\mathbb{R}),\ S^{-1}=\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix}\Bigg{\}},\end{split}$ (13)
Where $\mathrm{GL}_{2r}\,(\mathbb{R})$ denotes the set of all non-singular matrices in $\mathbb{R}^{2r\times 2r}$.
@@
In our experiments, we choose $S=\begin{bmatrix}\alpha I_{r}&0\\ 0&I_{r}\end{bmatrix}$, obtaining:
$\Delta W^{(0)}_{*}=\alpha U_{1,r}V^{\top}_{r,2r}\in\mathcal{M}_{r},\quad\alpha\in\mathbb{R}\setminus\{0\}.$ (14)
Surprisingly, the result of Theorem 4.1 is related to (Wang et al., 2024), where the authors did not consider the Riemannian setting and were not concerned with parametrization-free optimization.
Randomized SVD for efficient initialization.
@@
The computation of the $2r$-truncated SVD from Theorem 4.1 of the full loss gradient has asymptotical complexity $\mathcal{O}(\min\{m,n\}mn)$, which can represent significant computational demand for larger models. To overcome this problem, we propose to use a randomized SVD with power iterations (see (Halko et al., 2011)), as is described in the Algorithm 1. In a nutshell, we need to compute $\left(\nabla_{W}\mathcal{L}(W)\nabla_{W}\mathcal{L}(W)^{\top}\right)^{q}Y$, where $Y=\nabla_{W}\mathcal{L}(W)\Omega$ and $\Omega$ is sampled from standard normal distribution. This iteration can be done in a robust manner with the help of QR decompositions. The steps 2, 5, 7, and 10 correspond to matrix multiplications with gradients and are computed using the trick from (11). The larger the hyperparameters $p$ and $q$ are, the more accurate approximation we obtain. This method has asymptotic complexity of $\mathcal{O}((m+n)r^{2})$ and in addition $2(q+1)$ backward calls.
@@
2:  $Y:=\nabla_{A}\,\mathcal{L}(W+A\Omega^{\top})|_{A=0}=\nabla_{W}\,\mathcal{L}(W)\Omega.$
3:  $Y:=\mathtt{qr}(Y).\mathtt{Q}.$ // $\mathcal{O}\left(mr^{2}\right)$
4:  for $i:=1,\dots,q$ do
5:     $Y:=[\nabla_{B}\,\mathcal{L}(W+YB^{\top})|_{B=0}]^{\top}=\nabla_{W}\,\mathcal{L}(W)^{\top}Y.$
6:     $Y:=\mathtt{qr}(Y).\mathtt{Q}.$ // $\mathcal{O}\left(nr^{2}\right)$
7:     $Y:=\nabla_{A}\,\mathcal{L}(W+AY^{\top})|_{A=0}=\nabla_{W}\,\mathcal{L}(W)Y.$
8:     $Y:=\mathtt{qr}(Y).\mathtt{Q}.$ // $\mathcal{O}\left(mr^{2}\right)$
9:  end for
10:  $Q^{\top}\nabla_{W}\,\mathcal{L}(W):=\nabla_{B}\,\mathcal{L}(W+QB^{\top})^{\top}|_{B=0}.$
11:  $U,\Sigma,V^{\top}:=\mathtt{truncSVD}\left(Q^{\top}\nabla_{W}\,\mathcal{L}(W),r\right)$ // $\mathcal{O}\left(nr^{2}\right)$
12:  return  $QU,\Sigma,V^{\top}$ // $\mathcal{O}\left(mr^{2}\right)$
4.3 RiemannLoRA Optimization Algorithms
In Algorithm 3, we present the SGD with Heavy-Ball (Polyak, 1964) version of the RiemannLoRA fine-tuning algorithm. The first $3$ steps are dedicated to the computation of the initial guess for RiemannLoRA via BackPropRSVD for given layer weights and oversampling parameter $p$, that increases the accuracy of the singular approximation. The $4$-th step initializes Heavy-Ball momentum matrices and exponentially smoothed momentum norms. The part from step $5$ to step $18$ encapsulates the Riemannian optimization cycle on the fixed-rank manifold. First of all, in the $6$-th step, we calculate a representation of the manifold point with different orthogonalization (see the equation (4)) required for the subsequent calculation of the Riemannian gradient components at the current point in one backward call (steps $7,8$). Then the algorithm transports the Heavy-Ball tangent direction to the current point via the Algorithm 2 that provides a simple but effective implementation of (7). In the steps $9,10$, we compute the final optimization direction on the tangent space of the given point with a Heavy-Ball momentum coefficient $\beta$. Steps from $11$ to $14$ are dedicated to accumulating the momentum’s norm, similar to how it is done in the standard Adam. The step $15$ implements the SVD retraction step via vanilla truncated SVD (see (9)). The optimization cycle finalizes with saving the obtained minimization direction for momentum in the step $16$ and calculation of the new point representation in the step $17$. This approach has asymptotical complexity of $\mathcal{O}((m+n)r^{2}+r^{3})$ and $(2(q+1)+\mathtt{max\_iters})$ amount of backward calls.
Algorithm 2 VectorTransport
Require: Tangent vector $(A,B),A\in\mathbb{R}^{m\times 2r},B\in\mathbb{R}^{n\times 2r}$ from other tangent space in $2r$ representation (see (6)), current tangent vector foot point factors $A_{L},B_{R}$ with orthonormal columns.
Ensure: $(\dot{A},\dot{B})$ — parameters of orthoprojection of $(A,B)$ onto the current tangent space.
Function:
1:  return  $(A-A_{R}(A_{R}^{\top}A))(B^{\top}B_{R}),B(A^{\top}A_{L})$ // $\mathcal{O}\left((m+n)r^{2}\right)$
Algorithm 3 RiemannLoRA
@@ 4.3 RiemannLoRA Optimization Algorithms
Function:
1:  $A_{L},\_,B^{\top}:=\mathtt{BackPropRSVD}(2r,p,q,\mathcal{L},W)$. // $\mathcal{O}\left((m+n)r^{2}\right)$
2:  $A_{L},B^{\top}:=A_{L}[\colon,\colon r],\;B[\colon,r\colon]^{\top}$.
3:  $W^{\prime}:=W-A_{L}B^{\top}$. // $\mathcal{O}\left(mn\right)$
4:  $A_{\mathrm{HB}},B_{\mathrm{HB}},S_{A},S_{B}:=0$.
5:  for $i:=0,\dots,\mathtt{max\_iters}$ do
9:     $\dot{A}_{\mathrm{prev}},\dot{B}_{\mathrm{prev}}:=\mathtt{VectorTransport}\left((A_{\mathrm{HB}},B_{\mathrm{HB}}),A_{L},B_{R}\right).$ // $\mathcal{O}\left((m+n)r^{2}\right)$
11:     if $\mathtt{simulate\_Adam}$ then
14:     end if
15:     $U,\Sigma,V^{\top}:=\mathtt{truncSVD}\left(\left[\eta\dot{A},A_{L}\right],\left[B_{R},\eta\dot{B}+B\right]\right)$ // $\mathcal{O}\left((m+n)r^{2}+r^{3}\right)$
16:     $A_{\mathrm{HB}},B_{\mathrm{HB}}:=[\dot{A},A_{L}],\;[B_{R},\dot{B}]$
17:     $A_{L},B:=U,\Sigma V^{\top}$ // $\mathcal{O}\left(nr^{2}\right)$
18:  end for
19:  return  $A_{L},B$.
Also note that each step of the vanilla LoRA can be considered as moving in certain subparts of the tangent plane, i.e., along the vectors $(\dot{A},0)\in\mathcal{T}_{\Delta W}\mathcal{M}_{r}$ and $(0,\dot{B})\in\mathcal{T}_{\Delta W}\mathcal{M}_{r}$. In the RiemannLoRA we seek for the direction in the whole tangent space.

### v2

@@ 3.2 Riemannian optimization
When constructing the algorithms, we need to work with three key objects: Riemannian gradient, retraction and vector transport. The Euclidean gradient is a direction of the steepest local increase $F$. Therefore, it is common to use the Riemannian gradient — the direction of the steepest local increase of corresponding smooth function value along the manifold, which lies in the tangent space (Absil et al., 2009, chap. 3.6). Given the Euclidean gradient $\nabla F$, one may endow the tangent space $\mathcal{T}_{X}\mathcal{M}_{r}$ with a natural scalar product and derive a formula for the direction of the local steepest ascent of $F$ with respect to the manifold. This unique direction is called the Riemannian gradient and can be computed as follows:
$\mathrm{grad}\,F(X)=P_{\mathcal{T}_{X}\mathcal{M}_{r}}\,(\nabla F(X)),\quad X\in\mathcal{M}_{r}.$ (7)
A simple and robust retraction that maps a tangent step $\xi$ (for example, $\xi$ is a negative Riemannian gradient) back to the manifold is the truncated SVD:
$R_{X}(\xi)\equiv R(X+\xi)=\operatorname{SVD}_{r}(X+\xi),$ (8)
i.e., the best rank‑$r$ approximation of $X+\xi$ in Frobenius norm. Note that here we do not need to compute the full SVD and can utilize low-rank structure of $X$ and $\xi$, leading to $\mathcal{O}((m+n)r^{2}+r^{3})$ operations (Absil & Oseledets, 2015). Finally, because tangent spaces change from an optimization step to step, momentum (an accumulated tangent vector) must be moved between them via a vector transport. For embedded manifolds like $\mathcal{M}_{r}$, a standard choice is the projection transport
$\mathcal{T}_{Y\to X}(\xi)\;=\;P_{T_{X}\mathcal{M}_{r}}(\xi),\qquad\xi\in\mathcal{T}_{Y}\mathcal{M}_{r},$ (9)
which simply reprojects the same ambient matrix $\xi$ onto the new tangent space at $X$.
4 Riemannion
In this section, we focus on the setting of parameter-efficient fine-tuning and, hence, to the fixed-rank manifold. For fine-tuning of one layer the optimization problem becomes:
$\mathcal{L}(W+\Delta W)\to\min_{\Delta W\in\mathcal{M}_{r}},$
where $\mathcal{L}$ is a differentiable loss function. Note that optimizing on $\mathcal{M}_{r}$ removes the ambiguity of factorized parameterizations, because all computations are carried out in the intrinsic space of the product $X$ rather than in any particular factorization. So the formulas we write below will naturally be reparameterization-invariant. Let $G_{t}=P_{T_{W}\mathcal{M}_{r}}(\nabla\mathcal{L}(W_{t}))$ be the Riemannian gradient. A Riemannian heavy‑ball (Polyak, 1964) momentum step reads
$M_{t}\;=\;\beta\,\widehat{M}_{t-1}\;+\;G_{t},\quad\widehat{M}_{t-1}\;=\;\mathcal{T}_{W_{t-1}\to W_{t}}(M_{t-1}),$ (10)
$\Delta W_{t+1}\;=\;R\left(\Delta W_{t}-\eta\,M_{t}\right),$ (11)
with $M_{0}=0$, momentum parameter $\beta\in[0,1)$, and stepsize $\eta>0$.
Let us now discuss how to introduce a Muon-like variant of this iteration, which we refer to as Riemannion. A direct projection onto the set of orthogonal matrices $\mathrm{Ortho}(M_{t})$ presents two challenges. First, such a step does not respect the underlying Riemannian geometry. To address this, we propose to find the best approximation of $\mathrm{Ortho}(M_{t})$ on the tangent plane $T_{W}\mathcal{M}_{r}$. Such a solution is given via the projection onto the tangent plane $P_{T_{W}\mathcal{M}_{r}}(\mathrm{Ortho}(M_{t}))$. However, a second issue arises: although this projection is low-rank, its computation remains inefficient because the input matrix is of full rank. Let us notice that $M_{t}\in T_{\Delta W_{t}}\mathcal{M}_{r}$ and, hence, is of rank at most $2r$ (Section 3.2). At the same time, the LMO interpretation (Section 3.1) provides several admissible low-rank solutions, including one where $\mathrm{Ortho}r(\cdot)$ replaces only the first $2r$ singular values with $1$, while all others are set to 0. Consequently, we obtain the following update rule:
$\widetilde{M}_{t}=P_{T_{\Delta W_{t}}\mathcal{M}_{r}}(\mathrm{Ortho}_{r}(M_{t})).$ (12)
Note that $\mathrm{Ortho}_{r}(\cdot)$ exactly preserves the column and row spaces of $M_{t}\in T_{\Delta W_{t}}\mathcal{M}_{r}$. Although $P_{T_{\Delta W_{t}}\mathcal{M}_{r}}(\mathrm{Ortho}_{r}(M_{t}))$ does not yield singular values exactly equal to $1$, in practice they remain in a close proximity. In particular, in our experiments they were always in the interval $(0.9,1.1)$. This behavior is reminiscent of the Newton–Schulz iteration, which likewise produces approximate singular values. To eliminate this inexactness and obtain an accurate solution in the intersection of $T_{\Delta W_{t}}\mathcal{M}_{r}$ with the set of matrices whose first $2r$ singular values are exactly $1$, one could apply various strategies. For example, one may formally apply the alternating projection method (Cheney & Goldstein, 1959; Boyd & Dattorro, 2003, Theorem 4, Section 2):
$\widetilde{M}_{t}=P_{T_{\Delta W_{t}}\mathcal{M}_{r}}(\mathrm{Ortho}_{r}(\dots P_{T_{\Delta W_{t}}\mathcal{M}_{r}}(\mathrm{Ortho}_{r}(M_{t})))).$ (13)
As an alternative, we could solve the LMO problem on the tangent plane:
$\widetilde{M}_{t}\in\operatorname*{Arg\,max}_{\begin{subarray}{c}S\in T_{\Delta W_{t}}\mathcal{M}_{r}\\ \|S\|_{2}\leq 1\end{subarray}}\ \langle M_{t},S\rangle,$ (14)
where we proposed to apply observations from (Cesista, 2025) to the fixed-rank manifold. In Appendix D, we show that (12) is an approximate solution to the maximization problem (14). Applying (13) or accurately solving (14) on the intersection of two convex sets using convex optimization methods is more computationally expensive, and our experiments indicate that it has little to no impact on the overall convergence of the optimizer.
Algorithm 1 OrthoLR (efficient computation of $\mathrm{Ortho}_{r}(\xi)$ for $\xi\in T_{X}\mathcal{M}_{r}$).
Require: $\xi\in T_{X}\mathcal{M}_{r}$ given by $(\dot{A},\dot{B})$ from (5); $A_{L},B_{R}$ such that $X=A_{L}GB_{R}^{\top}$ as in (4)
Ensure: $A\in\mathbb{R}^{m\times 2r},B\in\mathbb{R}^{m\times 2r}$: $AB^{T}=\mathrm{Ortho}(\xi)$.
[…]
Algorithm 2 ProjectLR (efficient computation of $P_{T_{X}\mathcal{M}_{r}}(Z)$ for a rank-$r^{\prime}$ matrix $Z$).
Require: $A\in\mathbb{R}^{m\times r^{\prime}},B\in\mathbb{R}^{n\times r^{\prime}}$ such that $Z=AB^{\top}$; $A_{L},B_{R}$ such that $X=A_{L}GB_{R}^{\top}$ as in (4).
Ensure: $\xi=P_{T_{X}\mathcal{M}_{r}}(Z)$ given by $(\dot{A},\dot{B})$ from (5).
[…]
Let us finally show that $\widetilde{M}_{t}$ from (12) can be computed efficiently using $\mathcal{O}((m+n)r^{2}+r^{3})$ arithmetic operations. Indeed, first of all we need to apply $\mathrm{Ortho}_{r}(\cdot)$ to a tangent vector $M_{t}$. From (6), we know that a tangent vector can be represented in a form of a rank-$2r$ matrix. Such a representation can always be transformed into the compact SVD form with 2 QR decompositions and a single full SVD of a $2r\times 2r$ matrix, see Algorithm 1 called OrthoLR. As a next step, we need to project the obtained result (decomposed matrix of rank $2r$) onto the tangent plane. The operation can also be done efficiently via (6) and is summarized in Algorithm 2 called ProjectLR.
5 Locally optimal initialization (LOI)
Once the theoretical framework for the Riemannian optimizer is established, it is natural to consider an initialization scheme that accounts for the underlying Riemannian geometry. Given any $\Delta W\in\mathcal{M}_{r}$, we may write
$\mathcal{L}(W)=\mathcal{L}(\underbrace{W-\Delta W}_{W^{\prime}}+\Delta W)=\mathcal{L}\left(W^{\prime}+\Delta W\right).$ (15)
@@ 5 Locally optimal initialization (LOI)
[…]
@@
In our experiments, we use $S=\begin{bmatrix}\alpha I_{r}&0\\ 0&I_{r}\end{bmatrix}$, obtaining: $\Delta W^{(0)}_{*}=\alpha U_{1,r}V^{\top}_{r,2r}\in\mathcal{M}_{r},\alpha\in\mathbb{R}\setminus\{0\}$.
Interestingly, Theorem 5.1 relates to the findings of Wang et al. (2024), although their analysis neither adopts a Riemannian framework nor addresses parametrization-free optimization. Our optimizer further differs from Zhang & Pilanci (2024) and Mo et al. (2025, LORO) in that it avoids inversion of the Gram matrix. As a result, the method remains stable as $\|\Delta W^{(0)}_{\ast}\|\to 0$. Empirical results indicate that initializing $\Delta W^{(0)}_{\ast}$ with a small norm leads to improved performance. The procedure for selecting the scaling parameter $\alpha$ is described in Appendix F.
6 Single backward-pass gradient trick
This section introduces an efficient method for computing gradient-times–matrix in a matrix-free way, at a computational cost equivalent to a single backward pass. We then apply this technique in both the LOI initialization procedure in Algorithm 3 and the Riemannion optimizer in Algorithm 4.
The calculation of the full fine-tuning loss gradient $\nabla_{W}\mathcal{L}\left(W\right)$ with pretrained parameters $W\in\mathbb{R}^{m\times n}$ is computationally expensive. At the same time, for our framework we only need to compute the products $(\nabla_{W}\mathcal{L}\left(W\right)^{\top}M)$ or $(\nabla_{W}\mathcal{L}\left(W\right)N)$ for some $M\in\mathbb{R}^{m\times r},N\in\mathbb{R}^{n\times r}$. Using the trick from (Novikov et al., 2022), we may calculate both quantities simultaneously using a single forward-backward pass with a doubled rank representation.
First, we initialize differentiable parameters $Z_{1}=0\in\mathbb{R}^{m\times r}$ and $Z_{2}=0\in\mathbb{R}^{n\times r}$ and perform a simple forward pass $L:=\mathcal{L}\left(W+Z_{1}N^{\top}+MZ_{2}^{\top}\right)$. This step does not violate the pipelines of LoRA framework since it is equivalent to a standard LoRA forward pass with special adapter:
$Z_{1}N^{\top}+MZ_{2}^{\top}=\begin{bmatrix}Z_{1}&M\end{bmatrix}\begin{bmatrix}N&Z_{2}\end{bmatrix}^{\top}=\tilde{A}\tilde{B}^{\top},\quad\tilde{A}\in\mathbb{R}^{m\times 2r},\tilde{B}\in\mathbb{R}^{n\times 2r}.$
Then we invoke an autodiff algorithm for value $L$, which ensures us both:
$\begin{split}&\nabla_{Z_{1}}L=\nabla_{Z_{1}}\mathcal{L}\left(W+Z_{1}N^{\top}+MZ_{2}^{\top}\right)|_{Z_{1}=0,Z_{2}=0}=\nabla_{W}\mathcal{L}(W)N,\\ &\nabla_{Z_{2}}L=\nabla_{Z_{2}}\mathcal{L}\left(W+Z_{1}N^{\top}+MZ_{2}^{\top}\right)|_{Z_{1}=0,Z_{2}=0}=\nabla_{W}\mathcal{L}(W)^{\top}M.\end{split}$ (18)
Note, that if $\Delta W=A_{L}B^{\top}=AB_{R}^{\top}$ from $W+\Delta W$, then one may take $N=B_{R},M=A_{L},Z_{1}=0,Z_{2}=B,Y=W+\Delta W$ and the exact same operations work.
Notably, (18) is crucial for computation of the Riemannian gradient in (6). Therefore the proposed approach is a key building block that allows us to avoid forming the full fine-tune gradient loss and effectively compute matrix-matrix multiplications. This idea is the core for both: LOI initialization (Algorithm 3) and Riemannion optimizer (Algorithm 4).
Randomized SVD for efficient initialization
@@ 6 Single backward-pass gradient trick
The computation of the $2r$-truncated SVD of the full loss gradient, as required by Theorem 5.1, has asymptotic complexity $\mathcal{O}(\min\{m,n\}mn)$, which may be infeasible for large-scale models. In addition, explicitly forming this gradient is itself computationally expensive, and thus we need to avoid it in practice. To overcome this problem, we propose to use a randomized SVD with power iterations (see (Halko et al., 2011)), which we also enhance with our one-step gradient trick as is described in the Algorithm 3. In a nutshell, we need to compute $\left(\nabla_{W}\mathcal{L}(W)\nabla_{W}\mathcal{L}(W)^{\top}\right)^{q}Y$, where $Y=\nabla_{W}\mathcal{L}(W)\Omega$ and $\Omega$ is sampled from standard normal distribution. This iteration can be done in a robust manner using QR decompositions. The steps $2,4,5,6$ correspond to a trick from (18).
Overall, the LOI search procedure has asymptotic complexity $\mathcal{O}((m+n)r^{2})$, plus $2(q+1)$ additional backward passes. Moreover, since LOI search is executed only once before fine-tuning (which typically involves thousands of backward passes), its runtime overhead is negligible, taking merely $0.25\%$ of the total fine-tuning wall-clock time in our experiments. Section 7.1, about $2$ thousand optimization steps were carried out and Algorithm 3 accounted for merely $0.25\%$ of the total fine-tuning wall-clock time.
@@
[…]
3:for $i:=1,\dots,q$ do
4:  $Y:=\mathtt{qr}([\nabla_{B}\,\mathcal{L}(W+YB^{\top})|_{B=0}]^{\top}).\mathtt{Q}.$ // 11 backward pass $\mathcal{O}\left(nr^{2}\right)$
[…]
Efficient Riemannion implementation
The Riemannion optimizer is presented in Algorithm 4. Similar to vanilla Muon, it relies on a chosen Ortho procedure. Since the LoRA approach represents the fine-tuning shift $\Delta W$ with low-rank matrices, we employ an SVD-based Ortho procedure. For computational efficiency, this procedure is adapted to the fixed-rank manifold, as described in Section 4 and implemented in the OrthoLR (Algorithm 1). In detail, in step 1 we calculate the Riemannian gradient components via a single backward call (18). Then, in step $2$ the algorithm transports the Heavy-Ball tangent direction to the current point via Algorithm 2 that provides a simple but effective implementation of (6). In line $4$, we compute the final optimization direction on the tangent space of the given point with a Heavy-Ball momentum coefficient $\beta$. In line $5$ algorithm performs the retraction step.
The algorithm finalizes with saving the obtained minimization direction for momentum in line $6$ and calculating of a new point representation in step $7$.
Algorithm 4 One step of Riemannion
Require: Weight matrix $W^{\prime}\in\mathbb{R}^{m\times n}$, rank $r\in\mathbb{N}$, initial point $A_{L},B_{R}$, Heavy-ball momentum $A_{\mathrm{HB}},B_{\mathrm{HB}}$, step size $\eta$, momentum coefficient $\beta$, weight decay coefficient $\gamma$.
@@
[…]
Overall, one iteration of the Riemannion loop has asymptotic complexity $\mathcal{O}((m+n)r^{2}+r^{3})$ and additionally the same number of backward passes as vanilla LoRA. For comparison, the Euclidean Muon optimizer for LoRA (Section 3.1) exhibits the same asymptotic complexity. The complete framework and its time performance are summarized in Algorithm 5 and Appendix I.